OpenStudy (anonymous):

Jim picked a card from a standard deck, looked at it, and then put it back. He then picked a second card. What is the probability that Jim picked a heart or an ace on either pick? one over fifty two two over fifty two sixteen over fifty two seventeen over fifty two

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@Loser66 @arabpride

OpenStudy (anonymous):

@sourwing @Compassionate

OpenStudy (compassionate):

52 in a standard deck. Four aces in a standard deck Thirteen hearts in a standard deck 4/52 + 13/52 = 17/52

OpenStudy (anonymous):

can you help me answer 2 more questions @Compassionate

OpenStudy (compassionate):

Sorry, no... You caught me by luck... Try asking @Hero or @ranga

OpenStudy (anonymous):

@ranga please help me with these last two questions!

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

i posted the questions as a snapshot!

OpenStudy (ranga):

\[\sum_{n=3}^{12}20(0.5)^{n-1} = 20\sum_{n=3}^{12}(0.5)^{n-1} = 20(0.5^2 + 0.5^3+....+0.5^{11})\]Use sum of geometric series formula to add up the numbers in the parenthesis.

OpenStudy (ranga):

Sum of geometric series:\[a + ar + ar^2 + .... + ar^{n-1} = a\frac{1-r^n}{1-r}\\ \text{Here }a = 0.5^2;~~r = 0.5;~~n = 10\]

OpenStudy (anonymous):

I get 0.124755859

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

That is not what I am getting.

OpenStudy (ranga):

\[20(0.5^2 + 0.5^3+....+0.5^{11}) = 20 * 0.5^2 * \left(\frac{1 - 0.5^{10}}{1 - 0.5}\right) = ?\]

OpenStudy (ranga):

When using calculator be sure to make use of parenthesis.

OpenStudy (anonymous):

I got the answer! @ranga can you help me with the question below it!

OpenStudy (ranga):

What answer did you get?

OpenStudy (anonymous):

9.99

OpenStudy (ranga):

Yes!

OpenStudy (ranga):

For the last one, they want a sigma notation to sum term 4 through term 15. That gives a hint the sigma notation ought to be \(\sum_{n=4}^{15}\) and so it has got to be either 1 or 3. Try n = 1 in 1 and 3 and see which one gives the first term as 5 (because in the table when n = 1, \(a_n = a_1 = 5\).

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

would the correct answer be C @ranga

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@ganeshie8

OpenStudy (ranga):

First we eliminated the second and fourth choices because the sigma runs from n=1 to 15 in those choices whereas the problem states they want the sum from 4th term to 15th term and so the sigma should run from n = 4 to 15 which are the 1st and 3rd choices. In the first choice, the nth term is indicated by the expression after the sigma which is: \(5(-2)^{n-1}\). If you put n = 1, you get the first term to be: \(5(-2)^{1-1}= 5(-2)^0 = 5*1 = 5\) which is what the table indicates the first term is when n = 1. In the third choice, the nth term is: \(5(-2)^{n+1}\). If you put n = 1, you get the first term to be: \(5(-2)^{1+1}= 5(-2)^2 = 5*4 = 20\) which is NOT what the table indicates the first term is when n = 1.